Suppose you are committed to owning a $245,000 Ferrari. If you believe your mutual fund can achieve an annual rate of return of 11.2 percent and you want to buy the car in 9 years (on the day you turn 30), how much must you invest today? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

245000=P*(1.112)^9

P=245000/(1.112)^9

=$245000*0.384642186

which is equal to

=$94,237.34(Approx).